OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form and in slope-intercept form. Passing through (-3,-4) and parallel to the line whose equation is y=-3x+2 Write an equation for the line in point-slope form..... Write an equation for the line in slope-intercept form.....

OpenStudy (rulnick):

slope int form: y = -3x - 13

OpenStudy (rulnick):

point slope form: y-(-4) = -3(x-(-3)) ...

OpenStudy (rulnick):

or y+4 = -3(x+3).

OpenStudy (rulnick):

Note that the original line, y = -3x + 2, has slope -3, as do all lines parallel to it.